### v1

@@ 3.3.2 Environment sustainability versus safety, transparency, non-discrimination, and traceability
Transfer learning, particularly through parameter- and ration-based methods, holds the potential to cultivate sustainable AI systems while trading off non-discrimination and accuracy and ensuring fair transfer in the domain adaptation [193, 194]. This approach enhances the interpretability of models [195, 196], making it easier to understand how AI makes decisions. Furthermore, explainable AI plays a crucial role in assessing both the performance and validity of models post-transfer learning, providing insights into their operation [197]. However, challenges arise when the original, pre-trained model operates as a “black box” without transparency, necessitating potentially costly reprogramming efforts to enable transfer learning [198]. The comparative costs of this reprogramming for black-box models versus training models from scratch have yet to be thoroughly explored.
@@
Machine unlearning: Machine unlearning, as previously discussed, addresses privacy issues in machine learning by specifically eliminating individuals’ data from trained models, thereby improving non-discrimination and trust in AI systems. Unlike the resource-intensive process of developing a new model from scratch to adhere to the “right to be forgotten,” machine unlearning methods, such as exact and approximate unlearning, offer a more eco-friendly alternative by lowering computational costs. Exact unlearning, exemplified by techniques like checkpoint training [176] or SISA training with only the model trained with the subset containing to-be-forgotten data retrained [100], presents a favorable balance between unlearning accuracy and time. On the other hand, approximate unlearning employs the use of influence functions and a gradient-based deletion approach to diminish or restrict the impact of an individual’s data on the model, achieving “approximate” statistical non-identifiability [104, 103].
@@
Retrieval-augmented generation (RAG): presents a cutting-edge technique in natural language processing (NLP), merging the capabilities of pre-trained language models with information retrieval methods to elevate text creation [177]. This approach enables the model to dynamically access and integrate pertinent documents or data excerpts from a comprehensive corpus, infusing the text generation process with external information. The process involves using user queries to pinpoint relevant documents within a knowledge base, which are then combined with the query to form a prompt that supplies additional context for the response generation. In contrast to fine-tuning large language models (LLMs), RAG stands out as more environmentally sustainable due to its elimination of training expenses and the absence of costs associated with storing additional or modified LLMs.
@@
Model Compression: Reducing model size without compromising performance is essential for environmentally sustainable AI, as larger models result in higher inference times and increased energy consumption and memory storage. This survey investigates four techniques applied during the training and post-training phases: pruning, quantization, low-rank factorization, and knowledge distillation. Pruning eliminates redundant parameters, neurons, layers, and filters, particularly in image processing [178, 179]. Quantization and Low-factor factorization can be applied to convolutional and fully connected layers, while the former reduces the number of bits for each weight and the latter decomposes a large matrix into smaller matrices of parameters with proper factorization and rank selection [180, 181, 182, 183]. Knowledge distillation, focusing on classification-based tasks, replaces large layers with smaller ones through a teacher-student learning framework [184, 185, 186].
While model compression positively impacts environmental sustainability, it adversely affects fairness measures and the model’s traceability. Model compression can amplify existing biases in AI models, undermining non-discrimination efforts [199, 200, 201]. Additionally, it can be used to remove watermarks within the model, thereby reducing traceability [202]. The impact on transparency is multifaceted: though compression can decrease interpretability, hindering transparency [203], explainable AI techniques can be utilized to mitigate this effect [204, 205, 206]. Notably, compressed models can offer benefits in privacy protection [207].
@@ 3.4.1 Overview
The concept of traceability pertains to the ability to relate the unique identifiable entities in a verifiable way by all parties involved, from the development phase of an AI system to its use and deployment [208]. This concept is critical to the success of AI systems in terms of ensuring transparency, accountability, and governance [209].
-Regulations incorporating traceability: One significant regulation that includes traceability as a key component is the European Union’s proposed AI Act [21]. The AI Act emphasizes the importance of traceability for high-risk AI systems by requiring comprehensive documentation of the AI system’s development, deployment, and operational phases. This documentation is critical for understanding how high-risk AI systems are developed and how they perform throughout their lifetime, enabling the verification of compliance with the Act’s requirements, monitoring of operations, and post-market monitoring. The AI Act mandates that technical documentation must contain the detailed information necessary to assess the AI system’s compliance with relevant requirements, including the general characteristics, capabilities, limitations of the system, algorithms, data training, testing, and validation processes used, as well as documentation on the relevant risk management system. Furthermore, the AI Act requires high-risk AI systems to technically allow for automatic recording of events (i.e., logs) throughout the system’s lifetime, ensuring that activities related to the AI system can be traced back and verified.
In addition to the EU’s AI Act, the U.S. Government Accountability Office (GAO) has also outlined traceability as a critical aspect of responsible AI deployment [210]. The GAO’s AI Accountability Framework includes traceability as one of its principles, advising entities to document the results of monitoring activities and any corrective actions taken to promote traceability and transparency. This approach enhances accountability by ensuring that decisions and actions taken by AI systems can be traced back to their source, facilitating oversight and the identification of issues that may require remediation.
@@
-The role of traceability in AI regulation: The concept of traceability is crucial in AI regulation, providing a fundamental structure that guarantees AI systems are effectively monitored and scrutinized, held accountable, and conform to ethical norms and legal mandates [42]. By enabling the ability to track and understand all actions, decisions, and processes involved in the AI lifecycle, traceability fosters a higher degree of transparency. This is essential not only for the creators of AI systems but also for regulators and the public, ensuring that any decisions made by AI can be scrutinized and understood, thereby upholding accountability, especially when these decisions significantly impact individuals or society at large.
In the realm of AI regulations and guidelines, traceability is often a stipulated requirement. This entails maintaining comprehensive records of the data, models, and processes utilized throughout the development and deployment phases of AI systems. Such meticulous documentation simplifies the process of demonstrating compliance with both legal and ethical standards [42]. For instance, the EU AI Act emphasizes the importance of keeping detailed records as a means to facilitate adherence to regulatory frameworks, thereby underscoring the critical role of traceability in regulatory compliance [21].
Furthermore, traceability is integral to fostering ethical and responsible AI development and use [211]. It provides the necessary infrastructure to audit and review the behavior of AI systems thoroughly. This capability is crucial for identifying, analyzing, and rectifying biases, errors, or any unintended consequences that may emerge throughout an AI system’s lifecycle. Consequently, traceability supports the operationalization of ethical principles and the practice of responsible AI, ensuring that AI systems are developed and utilized in a manner that aligns with societal values and norms.
From a risk management perspective, traceability equips organizations with the tools to identify and mitigate potential issues early on [212]. It allows for the continuous monitoring and evaluation of AI systems to ensure their operation remains within the intended ethical and regulatory parameters. By enabling early detection and correction of problems, traceability significantly contributes to minimizing risks associated with AI systems, thereby ensuring their safe and effective deployment.
Lastly, traceability plays a crucial role in enhancing trust in AI technologies among users and the broader public. The knowledge that the workings and decisions of an AI system can be traced back to their origins fosters confidence in the technology’s reliability and fairness [213]. In an era where trust in AI is paramount for widespread adoption, traceability emerges as a key factor in building and maintaining this trust, thereby facilitating a more receptive environment for AI technologies.
@@
Papers Category Techniques Impact Discussion with the Rest Four Aspects in AI Regulation
[214], [215], [216] Blockchain Blockchain-based Federated Learning Recording each node’s training data and process for backtracking and audit Environmental sustainability (+), safety (+), transparency (+)
[217], [218] Water marking Direct embedding in data Enhancing source tracking and audit trails Safety (+), transparency (+)
[219], [220] Model watermarking Providing IP protection and regulatory compliance
[221], [222] Adversarial watermarking Proving ownership when AI models are subjected to various of attacks
[19] AI systems reproduc- eability Practices and tool support Using relevant tools, practices, and data models for traceability in their connection to building AI models and systems Safety (+), transparency (+), non-discrimination (+)
[223], [224] Open-source licenses Regulation Controlling the misuse and unauthorized use of open-source code Transparency (+)
@@ 3.4.2 Traceability versus safety, transparency, non-discrimination, and environmental sustainability
Blockchain: Blockchain technology employs its consensus mechanism and cryptographic technologies to enable users to record data on an immutable ledger, thus ensuring the integrity and traceability of the data. In the era of artificial intelligence, blockchain methods are increasingly used to store and share training data, model parameters, and training processes, enhancing traceability, privacy, accountability, and audibility [214].
@@
Functioning as a distributed ledger maintained by a peer-to-peer network, blockchain is often integrated into a federated learning (FL) framework. This approach decentralizes the model training process, distributing it across various nodes to maintain data privacy and security [225, 214, 215, 216]. In regulated industries such as healthcare and pharmaceuticals, where process reproducibility is crucial — including model reconstruction — blockchain provides a robust solution to meet these stringent requirements [214, 216]. More specifically, they deploy a multi-blockchain-based platform to create a comprehensive audit trail for all activities associated with the FL model training process [214]. In this setup, each node in the model uses a blockchain to store the training data and intermediate data derived from training sessions, each secured with hash values. Concurrently, a global blockchain, managed by a smart contract, orchestrates the overall training process. To minimize the storage burden on the blockchain, model parameters are kept in an InterPlanetary File System (IPFS) distributed network, with the ledger recording the corresponding IPFS-created addresses [215]. This approach significantly reduces the environmental impact of heavy data sharing and transmission in blockchain operations. Simultaneously, distributing the processing burden of the FL framework across multiple computational nodes minimizes both operational and computational overhead.
@@
Since blockchain techniques ensure not only traceability but also privacy on an immutable ledger, it is crucial for regulated companies like aerospace and pharmaceutical companies. These sectors require maintaining the highest quality standards for their products and the ability to reconstruct the entire production chain to gain crucial insights into the causes of any failures [214]. Therefore, it improves the model’s safety and transparency.
@@
Watermarking: In the context of AI that involves embedding a unique, invisible marker or pattern into AI-generated outputs (such as images, videos, or text) or within the AI model itself. This marker is designed to be robust against modifications and should be retrievable even after the data undergoes transformations or compression [226].
@@
Watermarks are subtly incorporated directly into content, such as images or video frames, through a process known as direct embedding [217, 218]. This method ensures the watermarks are both imperceptible to users and recoverable when needed. It is a technique frequently utilized in various forms of media and official documents. Model watermarking, on the other hand, involves adjusting a neural network’s internal parameters-like specific weights or biases-to embed a watermark [219]. This is done in such a way that the model’s performance remains unaffected [220]. The presence of the watermark can be confirmed through specific queries that prompt the model to produce pre-defined responses, revealing the watermark. Adversarial watermarking represents a more advanced technique [221, 222]. In this method, the process of embedding a watermark involves the creation of adversarial example inputs specifically designed to challenge the model. These inputs are subtly modified to include the watermark and are used during the model’s training phase. The model is thus trained to recognize these examples and respond in a specific manner, embedding the watermark more thoroughly within its operational framework. This type of watermarking embeds it deeply within the model’s decision-making patterns, making it particularly robust.
@@
Several tools have been developed to achieve reproducibility, which, in turn, supports traceability [19]. ModelDB [227] is an open-source system designed for versioning machine learning models, allowing users to index, track, and store modeling artifacts for later reproduction, sharing, and analysis. It emphasizes experiment tracking and provides a web-based interface for the visual representation and analysis of models. Code Ocean  [^17] https://codeocean.com/, Whole Tale  [^18] https://wholetale.org/, and The Renku Project  [^19] https://datascience.ch/renku/ are among the numerous online tools aimed at enhancing reproducibility. These platforms leverage cloud storage and containerization technologies, such as Docker, to capture the research environment fully. This enables the reuse, sharing, and reproduction of the entire research process. Code Ocean merges leading tools, languages, and environments to offer an end-to-end workflow focused on reproducibility. Whole Tale is a free, open-source platform that captures data, code, and the complete software environment, aspiring to redefine how computational and data-driven science is conducted and reproduced. The Renku Project combines a web platform with a command-line interface to support reproducibility, reusability, and collaboration. A broader set of tools supporting “methods reproducibility” research includes ZenML  [^20] https://zenml.io/, Binder  [^21] https://mybinder.org/, DVC (i.e., Data Version Control)  [^22] https://dvc.org/, Taverna  [^23] https://taverna.incubator.apache.org/, Kepler  [^24] https://kepler-project.org/, and VisTrails  [^25] https://www.vistrails.org/. These tools are designed to manage different aspects of the computational research lifecycle, including environment setup, code execution, data management, and the tracking of provenance and metadata. These tools provide essential capabilities for ensuring that AI research and applications are reproducible, which in turn supports the traceability of AI systems by documenting and validating the steps, data, and outcomes involved in model development and deployment.
@@
Open-source license: An open-source license is a type of license for computer software and other products that allows the source code to be used, modified, and distributed by anyone. Open-source licenses are designed to encourage collaboration and sharing, promoting the development of the software in a community-driven manner. These licenses allow the software to be freely used, modified, and shared under defined terms and conditions [228].
@@
Open-source licenses require that the source code be made available to the public. This transparency allows researchers, developers, and users to examine the algorithms, data processing methods, and decision-making processes within AI systems [229]. It helps in understanding how these systems work, identifying potential biases, errors, or vulnerabilities, and ensuring that the AI behaves as intended. By allowing anyone to access, modify, and distribute the source code, open-source licenses foster a collaborative environment. Open-source AI projects also enable other researchers to reproduce and verify the results claimed by the original developers. This is a fundamental aspect of scientific research that ensures the reliability and validity of AI technologies. Lastly, open-source projects typically maintain extensive documentation and use version control systems. This practice enhances traceability by providing a detailed history of changes, updates, and modifications made to the AI system over time.

### v2

@@ 3.3.2 Environment sustainability versus safety, transparency, non-discrimination, and traceability
Transfer learning, particularly through parameter- and ration-based methods, holds the potential to cultivate sustainable AI systems while trading off non-discrimination and accuracy and ensuring fair transfer in the domain adaptation (?, ?). This approach enhances the interpretability of models (?, ?), making it easier to understand how AI makes decisions. Furthermore, explainable AI plays a crucial role in assessing both the performance and validity of models post-transfer learning, providing insights into their operation (?). However, challenges arise when the original, pre-trained model operates as a “black box” without transparency, necessitating potentially costly reprogramming efforts to enable transfer learning (?). The comparative costs of this reprogramming for black-box models versus training models from scratch have yet to be thoroughly explored.
@@
Machine unlearning: Machine unlearning, as previously discussed, addresses privacy issues in machine learning by specifically eliminating individuals’ data from trained models, thereby improving non-discrimination and trust in AI systems. Unlike the resource-intensive process of developing a new model from scratch to adhere to the “right to be forgotten,” machine unlearning methods, such as exact and approximate unlearning, offer a more eco-friendly alternative by lowering computational costs. Exact unlearning, exemplified by techniques like checkpoint training (?) or SISA training with only the model trained with the subset containing to-be-forgotten data retrained (?), presents a favorable balance between unlearning accuracy and time. On the other hand, approximate unlearning employs the use of influence functions and a gradient-based deletion approach to diminish or restrict the impact of an individual’s data on the model, achieving “approximate” statistical non-identifiability (?, ?).
@@
Retrieval-augmented generation (RAG): presents a cutting-edge technique in natural language processing (NLP), merging the capabilities of pre-trained language models with information retrieval methods to elevate text creation (?). This approach enables the model to dynamically access and integrate pertinent documents or data excerpts from a comprehensive corpus, infusing the text generation process with external information. The process involves using user queries to pinpoint relevant documents within a knowledge base, which are then combined with the query to form a prompt that supplies additional context for the response generation. In contrast to fine-tuning large language models (LLMs), RAG stands out as more environmentally sustainable due to its elimination of training expenses and the absence of costs associated with storing additional or modified LLMs.
@@
Model Compression: Reducing model size without compromising performance is essential for environmentally sustainable AI, as larger models result in higher inference times and increased energy consumption and memory storage. This survey investigates four techniques applied during the training and post-training phases: pruning, quantization, low-rank factorization, and knowledge distillation. Pruning eliminates redundant parameters, neurons, layers, and filters, particularly in image processing (?, ?). Quantization and Low-factor factorization can be applied to convolutional and fully connected layers, while the former reduces the number of bits for each weight and the latter decomposes a large matrix into smaller matrices of parameters with proper factorization and rank selection (?, ?, ?, ?). Knowledge distillation, focusing on classification-based tasks, replaces large layers with smaller ones through a teacher-student learning framework (?, ?, ?).
While model compression positively impacts environmental sustainability, it adversely affects fairness measures and the model’s traceability. Model compression can amplify existing biases in AI models, undermining non-discrimination efforts (?, ?, ?). Additionally, it can be used to remove watermarks within the model, thereby reducing traceability (?). The impact on transparency is multifaceted: though compression can decrease interpretability, hindering transparency (?), explainable AI techniques can be utilized to mitigate this effect (?, ?, ?). Notably, compressed models can offer benefits in privacy protection (?).
@@ 3.4.1 Overview
The concept of traceability pertains to the ability to relate the unique identifiable entities in a verifiable way by all parties involved, from the development phase of an AI system to its use and deployment (?). This concept is critical to the success of AI systems in terms of ensuring transparency, accountability, and governance (?).
-Regulations incorporating traceability: One significant regulation that includes traceability as a key component is the European Union’s proposed AI Act (?). The AI Act emphasizes the importance of traceability for high-risk AI systems by requiring comprehensive documentation of the AI system’s development, deployment, and operational phases. This documentation is critical for understanding how high-risk AI systems are developed and how they perform throughout their lifetime, enabling the verification of compliance with the Act’s requirements, monitoring of operations, and post-market monitoring. The AI Act mandates that technical documentation must contain the detailed information necessary to assess the AI system’s compliance with relevant requirements, including the general characteristics, capabilities, limitations of the system, algorithms, data training, testing, and validation processes used, as well as documentation on the relevant risk management system. Furthermore, the AI Act requires high-risk AI systems to technically allow for automatic recording of events (i.e., logs) throughout the system’s lifetime, ensuring that activities related to the AI system can be traced back and verified.
In addition to the EU’s AI Act, the U.S. Government Accountability Office (GAO) has also outlined traceability as a critical aspect of responsible AI deployment (?). The GAO’s AI Accountability Framework includes traceability as one of its principles, advising entities to document the results of monitoring activities and any corrective actions taken to promote traceability and transparency. This approach enhances accountability by ensuring that decisions and actions taken by AI systems can be traced back to their source, facilitating oversight and the identification of issues that may require remediation.
@@
-The role of traceability in AI regulation: The concept of traceability is crucial in AI regulation, providing a fundamental structure that guarantees AI systems are effectively monitored and scrutinized, held accountable, and conform to ethical norms and legal mandates (?). By enabling the ability to track and understand all actions, decisions, and processes involved in the AI lifecycle, traceability fosters a higher degree of transparency. This is essential not only for the creators of AI systems but also for regulators and the public, ensuring that any decisions made by AI can be scrutinized and understood, thereby upholding accountability, especially when these decisions significantly impact individuals or society at large.
In the realm of AI regulations and guidelines, traceability is often a stipulated requirement. This entails maintaining comprehensive records of the data, models, and processes utilized throughout the development and deployment phases of AI systems. Such meticulous documentation simplifies the process of demonstrating compliance with both legal and ethical standards (?). For instance, the EU AI Act emphasizes the importance of keeping detailed records as a means to facilitate adherence to regulatory frameworks, thereby underscoring the critical role of traceability in regulatory compliance (?).
Furthermore, traceability is integral to fostering ethical and responsible AI development and use (?). It provides the necessary infrastructure to audit and review the behavior of AI systems thoroughly. This capability is crucial for identifying, analyzing, and rectifying biases, errors, or any unintended consequences that may emerge throughout an AI system’s lifecycle. Consequently, traceability supports the operationalization of ethical principles and the practice of responsible AI, ensuring that AI systems are developed and utilized in a manner that aligns with societal values and norms.
From a risk management perspective, traceability equips organizations with the tools to identify and mitigate potential issues early on (?). It allows for the continuous monitoring and evaluation of AI systems to ensure their operation remains within the intended ethical and regulatory parameters. By enabling early detection and correction of problems, traceability significantly contributes to minimizing risks associated with AI systems, thereby ensuring their safe and effective deployment.
Lastly, traceability plays a crucial role in enhancing trust in AI technologies among users and the broader public. The knowledge that the workings and decisions of an AI system can be traced back to their origins fosters confidence in the technology’s reliability and fairness (?). In an era where trust in AI is paramount for widespread adoption, traceability emerges as a key factor in building and maintaining this trust, thereby facilitating a more receptive environment for AI technologies.
@@
Papers Category Techniques Impact Discussion with the Rest Aspects in AI Regulation
(?, ?, ?) Blockchain Blockchain-based Federated Learning Recording each node’s training data and process for backtracking and audit Environmental sustainability (+), safety (+), transparency (+)
(?, ?) Water marking Direct embedding in data Enhancing source tracking and audit trails Safety (+), transparency (+)
(?, ?) Model watermarking Providing IP protection and regulatory compliance
(?, ?) Adversarial watermarking Proving ownership when AI models are subjected to various attacks
(?) AI systems reproduceability Practices and tool support Using relevant tools, practices, and data models for traceability in their connection to building AI models and systems Safety (+), transparency (+), non-discrimination (+)
(?, ?) Open-source licenses Regulation Controlling the misuse and unauthorized use of open-source code Transparency (+)
@@ 3.4.2 Traceability versus safety, transparency, non-discrimination, and environmental sustainability
Blockchain: Blockchain technology employs its consensus mechanism and cryptographic technologies to enable users to record data on an immutable ledger, thus ensuring the integrity and traceability of the data. In the era of artificial intelligence, blockchain methods are increasingly used to store and share training data, model parameters, and training processes, enhancing traceability, privacy, accountability, and audibility (?).
@@
Functioning as a distributed ledger maintained by a peer-to-peer network, blockchain is often integrated into a federated learning (FL) framework. This approach decentralizes the model training process, distributing it across various nodes to maintain data privacy and security (?, ?, ?, ?). In regulated industries such as healthcare and pharmaceuticals, where process reproducibility is crucial — including model reconstruction — blockchain provides a robust solution to meet these stringent requirements (?, ?). More specifically, they deploy a multi-blockchain-based platform to create a comprehensive audit trail for all activities associated with the FL model training process (?). In this setup, each node in the model uses a blockchain to store the training data and intermediate data derived from training sessions, each secured with hash values. Concurrently, a global blockchain, managed by a smart contract, orchestrates the overall training process. To minimize the storage burden on the blockchain, model parameters are kept in an InterPlanetary File System (IPFS) distributed network, with the ledger recording the corresponding IPFS-created addresses (?). This approach significantly reduces the environmental impact of heavy data sharing and transmission in blockchain operations. Simultaneously, distributing the processing burden of the FL framework across multiple computational nodes minimizes both operational and computational overhead.
@@
Since blockchain techniques ensure not only traceability but also privacy on an immutable ledger, it is crucial for regulated companies like aerospace and pharmaceutical companies. These sectors require maintaining the highest quality standards for their products and the ability to reconstruct the entire production chain to gain crucial insights into the causes of any failures (?). Therefore, it improves the model’s safety and transparency.
@@
Watermarking: In the context of AI that involves embedding a unique, invisible marker or pattern into AI-generated outputs (such as images, videos, or text) or within the AI model itself. This marker is designed to be robust against modifications and should be retrievable even after the data undergoes transformations or compression (?).
@@
Watermarks are subtly incorporated directly into content, such as images or video frames, through a process known as direct embedding (?, ?). This method ensures the watermarks are both imperceptible to users and recoverable when needed. It is a technique frequently utilized in various forms of media and official documents. Model watermarking, on the other hand, involves adjusting a neural network’s internal parameters-like specific weights or biases-to embed a watermark (?). This is done in such a way that the model’s performance remains unaffected (?). The presence of the watermark can be confirmed through specific queries that prompt the model to produce pre-defined responses, revealing the watermark. Adversarial watermarking represents a more advanced technique (?, ?). In this method, the process of embedding a watermark involves the creation of adversarial example inputs specifically designed to challenge the model. These inputs are subtly modified to include the watermark and are used during the model’s training phase. The model is thus trained to recognize these examples and respond in a specific manner, embedding the watermark more thoroughly within its operational framework. This type of watermarking embeds it deeply within the model’s decision-making patterns, making it particularly robust.
@@
Several tools have been developed to achieve reproducibility, which, in turn, supports traceability (?). ModelDB (?) is an open-source system designed for versioning machine learning models, allowing users to index, track, and store modeling artifacts for later reproduction, sharing, and analysis. It emphasizes experiment tracking and provides a web-based interface for the visual representation and analysis of models. Code Ocean  [^17] https://codeocean.com/, Whole Tale  [^18] https://wholetale.org/, and The Renku Project  [^19] https://datascience.ch/renku/ are among the numerous online tools aimed at enhancing reproducibility. These platforms leverage cloud storage and containerization technologies, such as Docker, to capture the research environment fully. This enables the reuse, sharing, and reproduction of the entire research process. Code Ocean merges leading tools, languages, and environments to offer an end-to-end workflow focused on reproducibility. Whole Tale is a free, open-source platform that captures data, code, and the complete software environment, aspiring to redefine how computational and data-driven science is conducted and reproduced. The Renku Project combines a web platform with a command-line interface to support reproducibility, reusability, and collaboration. A broader set of tools supporting “methods reproducibility” research includes ZenML  [^20] https://zenml.io/, Binder  [^21] https://mybinder.org/, DVC (i.e., Data Version Control)  [^22] https://dvc.org/, Taverna  [^23] https://taverna.incubator.apache.org/, Kepler  [^24] https://kepler-project.org/, and VisTrails  [^25] https://www.vistrails.org/. These tools are designed to manage different aspects of the computational research lifecycle, including environment setup, code execution, data management, and the tracking of provenance and metadata. These tools provide essential capabilities for ensuring that AI research and applications are reproducible, which in turn supports the traceability of AI systems by documenting and validating the steps, data, and outcomes involved in model development and deployment.
@@
Open-source license: An open-source license is a type of license for computer software and other products that allows the source code to be used, modified, and distributed by anyone. Open-source licenses are designed to encourage collaboration and sharing, promoting the development of the software in a community-driven manner. These licenses allow the software to be freely used, modified, and shared under defined terms and conditions (?).
@@
Open-source licenses require that the source code be made available to the public. This transparency allows researchers, developers, and users to examine the algorithms, data processing methods, and decision-making processes within AI systems (?). It helps in understanding how these systems work, identifying potential biases, errors, or vulnerabilities, and ensuring that the AI behaves as intended. By allowing anyone to access, modify, and distribute the source code, open-source licenses foster a collaborative environment. Open-source AI projects also enable other researchers to reproduce and verify the results claimed by the original developers. This is a fundamental aspect of scientific research that ensures the reliability and validity of AI technologies. Lastly, open-source projects typically maintain extensive documentation and use version control systems. This practice enhances traceability by providing a detailed history of changes, updates, and modifications made to the AI system over time.